Answer the same questions as in the above problem if:

TC = 120 + 50Q – 10Q2 + Q3



1
Expert's answer
2021-05-04T11:37:03-0400


As full question is not given , but these types of question require basic , equations

That are TVC,AVC and MC , after these equations further can be solved easily , so

As follows



The equations of the TVC, AVC, and MC functions.


 Calculating TVC


From the TC function, $100 is not affected by output (Q), and hence FC = $100. The remaining part of the function affected by output(Q) is the TVC function.


Thus, "TVC = 50Q-10Q^2+Q3"



Calculating AVC


"AVC = \dfrac {TVC}{Q}"


"= \dfrac {50Q-10Q^2+Q^3}{Q}"


"= \dfrac {Q(50-10Q+Q^2)}{Q}"


"= 50-10Q+Q^2"



Calculating MC


"MC = \dfrac {d}{dQ}(TC)"


"= \dfrac {d}{dQ}(120 + 50Q - 10Q^2 + Q^3)"

"= 50 - 20Q + 3Q^2"